# Prediction of causal genes at GWAS loci with pleiotropic gene regulatory effects using sets of correlated instrumental variables

### v2

@@ Introduction
For the diagrams with a single exposure (Fig. 1A,D), we set the true causal effect $c_{X}=0.3$. In the case of one instrument (Fig. 1A) we varied the instrument strength $a$ between $0.1$ (weak instrument), $0.5$ (mediocre instrument), and $0.8$ (strong instrument). In the case of multiple instruments (Fig. 1D), we simulated ten instruments with randomly sampled instrument strengths, either all strong (effect sizes greater than $0.5$) or all weak (effect sizes smaller than $0.03$).
@@
To test the effect of instrument correlation we simulated data with two instruments (Fig. 1B) where the pairwise correlation of the instruments varied between $0.01$, $0.3$, $0.7$, $0.9$ and $0.95$. The matrix $A=[a_{ij}]$ of instrument effect sizes, where $a_{ij}$ is the direct causal effect of $E_{i}$ on $X_{j}$, was generated randomly such that its determinant was greater than $0.5$ and there were no weak instruments ($a_{ij}$ greater than 0.6).
To test the effect of weak instruments, we simulated data with two instruments (Fig. 1B) with randomly generated non-zero pairwise correlation of the instruments and randomly generated matrix of instrument effect sizes $A$ with determinant either less than $0.2$ and effect size of each instrument less than $0.03$ or determinant greater than $0.5$ and effect size of each instrument greater than $0.5$.
To test robustness of the causal effect estimates, we simulated data with two instruments and two exposures (Fig. 1B), with randomly generated non-zero pairwise correlation between the instruments, randomly generated matrix of instrument effect sizes $A$ with determinant greater than $0.5$, and true causal effects of the exposures on the outcome equal to $c_{X_{1}}=0.2$ and $c_{X_{2}}=0.6$, but estimated causal effects for each exposure separately under the (false) hypothesis that the data was generated from Fig. 1A.
@@
We assumed that a randomly selected subset of at least $K$ eQTLs were causal and all others were associated with the exposures only through LD. For the causal eQTLs, we sampled the vector $\mathbf{a}$ of instrument effect sizes randomly between $0.7$ and $1.0$,
@@
For the simulations on the accuracy of the LD matrix, we generated datasets for the SLC22A3-LPA-PLG locus (centered on chr $6:161089307$) in liver as exposures. Causal effects of the genes on the outcome variable $Y$ were calculated to be $c_{\textit{SLC22A3}}=0.15$, $c_{\textit{LPA}}=-0.05$ and $c_{\textit{PLG}}=-0.27$ when seven instruments were used, and the same values were also used in the other simulations. We compared the situations where $L=3$ (LD threshold of 0.25), $L=4$ (LD threshold of 0.3), $L=5$ (LD threshold of 0.4), $L=6$ (LD threshold of 0.5), and $L=7$ instruments (LD threshold of 0.8) were used. We performed one-sample simulations where the sample size varied between $500-30000$. For the instrument effect sizes, we assumed that three eQTLs were causal for all genes (their effect sizes sampled from a uniform distribution within the range $0.7-1.0$), while additional eQTLs were assumed to be associated with the exposures only through LD.
For the simulations on weak instrument bias, we generated datasets within the same locus and setup as the LD accuracy simulations except here we kept $L=8$ eQTLs (LD threshold of 0.96) and varied the sample size between $500-10000$. For the instrument effect sizes, we again assumed that only three eQTLs were causal for all genes, their effect sizes sampled from a uniform distribution within the range $0.7-1.0$ for strong instruments and within the range $0.001-0.03$ for weak instruments. We confirmed that these instruments were indeed weak/strong by employing the conditional F-statistics computed using the Mendelian Randomization package [30] as well as the MVMR package [31] (S3 Fig).
@@
To simulate the effect of horizontal pleiotropy, we generated datasets for the SLC22A3-LPA-PLG locus with causal effects of the genes on the outcome variable $Y$ set to $c_{\textit{LPA}}=-0.05$ and $c_{\textit{PLG}}=-0.27$ and $c_{\textit{SLC22A3}}\in[0.0,0.4]$. We used eight instruments where only three eQTLs were causal for all genes, their effect sizes sampled from a uniform distribution within the range $0.7-1.0$. We estimated the effects in a misspecified estimation model where we performed MVMR only using LPA-PLG genes, assuming horizontal pleiotropy through one unknown/unmeasured gene SLC22A3 in the same locus.
@@
We compared the least squares (LS) (eq. (7)) and GMM (eq. (10)) estimators to three other multi-variable MR methods: Transcriptome-Wide MR (TWMR) [19], the mv_multiple estimator from TwoSampleMR package [32], and the mv_mvivw estimator from the Mendelian Randomization package [30].
@@
We used eQTL summary statistics from the Stockholm-Tartu Atherosclerosis Reverse Network Engineering Task (STARNET) study, a genetics-of-gene expression study of tissue samples from blood ($n=481$), atherosclerotic-lesion-free internal mammary artery (MAM, $n=552$), atherosclerotic aortic root (AOR, $n=539$), subcutaneous fat (SF, $n=573$), visceral abdominal fat (VAF, $n=531$), skeletal muscle (SKLM, $n=534$), and liver (LIV, $n=576$) obtained during open thorax surgery of 600 coronary artery disease (CAD) patients [26]. eQTL summary statistics from matching tissues in GTEx [12] were used for replication analyses.
@@
To identify lead SNPs for genome-wide significant GWAS loci with regulatory effects in a given tissue, we first created tissue-specific outcome data with GWAS summary statistics for all SNPs with non-zero (FDR $<5\%$) eQTL effect in the STARNET summary statistics for the tissue of interest. We filtered these eQTLs by their p-values, retaining only those with p-values below $5\times 10^{-8}$ in the GWAS study.
We iterated through these eQTLs, sorted by GWAS p-value, such that for the first iteration, the first eQTL would be the first lead SNP. We collected all eQTLs within a $0.5Mb$ radius and with non-zero LD with the lead SNP, and then removed the lead SNP as well as its collection of linked eQTLs from the list, repeating the procedure with the next lead SNP. We computed conditional F-statistics using the Mendelian Randomization [30] and MVMR [31] packages to verify instrument strength.
For each lead SNP and its collection of linked eQTLs, we first removed SNPs in perfect LD with the lead SNP and then treated the remaining eQTLs as instrumental variables in one causal diagram. We gathered information on genes associated with these eQTLs either in a specific tissue, or across all tissues, depending on the application, to complete the diagram.
@@ Results
Fig 2: Causal effect estimation from data simulating instrument correlations, instrument strength, and horizontal pleiotropy. (A) Causal diagram for the simulation of two exposures $X_{1}$ and $X_{2}$ for an outcome $Y$, with two shared instruments $E_{1}$ and $E_{2}$ with pairwise covariance $\alpha$ and matrix of instrument effect sizes $A=[a_{ij}]$. (B) Difference ($\hat{c}-c$) between estimated and true causal effect of $X_{1}$ (true effect size $c=0.2$) on $Y$ across 1,000 independently simulated datasets for a range of sample sizes. (C) Variance explained by the first two PCs for instrument covariances $\alpha=0.01$, $0.3$, $0.7$, $0.9$ and $0.95$. (D,E) Distribution of the difference between estimated and true causal effects for $X_{1}$ (left; true effect size $c=0.2$) and $X_{2}$ (right; true effect size $c=0.6$) across 1,000 independently simulated datasets and a range of sample sizes comparing strong $0.7-1.0$ vs weak $0.001-0.03$instruments; insets show a zoom along y-axis ranges. (F) False causal diagram used for inference where one causal exposure is missing. (G-H) Difference between estimated and true causal effect of $LPA$ (left, true effect size $c_{\textit{LPA}}=-0.05$) and $PLG$ (right, true effect size $c_{\textit{PLG}}=-0.27$) on $Y$ across 1,000 independently simulated datasets with sample size 2000 across a range of effect sizes of the hidden gene SLC22A3 mediating horizontal pleiotropic effects of the instruments on $Y$, showing both the estimates with the correct model (green) and the estimates with the hidden exposure model (orange). See Methods for simulation details.
@@
Fig 3: Causal effect estimation from data simulating LD matrix sampling bias. Each row shows a causal diagram for the simulation of overdetermined systems where the number of causal variants was equal to the number of exposures (first row, three exposures; second row, two exposures) and other variants were only correlated with the exposures and outcome by LD. (A, B, C) Comparison between the GMM estimator using an unbiased (B) and biased (C) LD matrix for the causal effect of simulated SLC22A3, for different LD pruning thresholds (corresponding to models with 3, 4, 5, 6 and 7 instruments). (D, E, F) Comparison of the least-squares, GMM, TWMR, MV multiple (TwoSampleMR), and MVIVW (Mendelian Randomization package) estimators using an unbiased (E) and biased (F) LD matrix for the causal effect of simulated MRAS and ESTY3. All dsitributions plots show the difference between estimated and true causal effects across 1,000 independently simulated datasets from the causal diagram on the correspinding row. See Methods for simulation details.
@@ Discussion
Extensive simulations confirmed the validity and usefulness of these theoretical results. Most surprisingly perhaps was the finding that the variance in causal effect estimates remained small, even at high correlation values between the instruments. Moreover we saw no benefit in using the computationally more expensive two-stage least squares estimator instead of the simpler least squares estimator – both were unbiased and had identical variance in all our simulations. As expected, horizontal pleiotropy or misspecified LD matrices introduce bias in the estimated causal effects, as in all applications of MR, but using conservative simulation parameters, we observed convergence to the correct estimates at sample sizes at the upper end of what is currently available.
@@
The fact that regulatory pleiotropy can extend over multiple cell types and tissues implies that MVMR must be run strictly speaking by including all gene-tissue pairs with cis-eQTL associations in a given GWAS locus as exposures in a single, large causal modelin order to limit the risk of horizonal pleiotropy. However, even though instrument sets can be highly correlated and the STARNET data contains only seven tissues, it was almost never possible to run such large models, as the requirement on the number of causal variants included in or tagged by the instrument set still stands. Moreover, data from some potentially relevant cell types or tissues will almost always be missing. To overcome these limitations, we see two ways forward.
@@ Acknowledgments
Conditional F-statistic estimation on simulated data of over-determined systems. (A) Distribution of conditional F-statistic using the Mendelian Randomization package (B) Distribution of conditional F-statistic using the MVMR package (C) Distribution of estimated causal effects for $ADAMTS7$ (true effect size $0.15$) for the estimator GMM, showing distributions across 2,000 independently simulated datasets across a range of sample sizes using discrete instruments with randomly generated covariances with real LD values from the ADAMTS7-CTSH locus on $15:79124475$ in the Aorta tissue. (D) Distribution of estimated causal effects for $ADAMTS7$ (true effect size $0.15$) for the estimator mvivw estimator using the Mendelian Randomization package which returns the estimates along with the conditional F-statistics.

### v3

@@ Introduction
For the diagrams with a single exposure (Fig. 1A,D), we set the true causal effect $c_{X}=0.3$. In the case of one instrument (Fig. 1A) we varied the instrument strength $a$ between $0.01$ (weak instrument), $0.3$ (mediocre instrument), and $0.8$ (strong instrument). In the case of multiple instruments (Fig. 1D), we simulated ten instruments with randomly sampled instrument strengths, either all strong (effect sizes between $0.1-0.3$) or all weak (effect sizes between $0.001-0.03$).
@@
To test the effect of instrument correlation we simulated data with two instruments (Fig. 1B) where the pairwise correlation of the instruments varied between $0.01$, $0.3$, $0.7$, $0.9$ and $0.95$. The matrix $A=[a_{ij}]$ of instrument effect sizes, where $a_{ij}$ is the direct causal effect of $E_{i}$ on $X_{j}$, was generated randomly such that its determinant was greater than $0.05$ and there were no weak instruments ($a_{ij}$ between $0.1-0.3$).
To test the effect of weak instruments, we simulated data with two instruments (Fig. 1B) with randomly generated non-zero pairwise correlation of the instruments and randomly generated matrix of instrument effect sizes $A$ with determinant either less than $0.001$ and effect size of each instrument less than $0.001$ or determinant greater than $0.05$ and effect size of each instrument between $0.1-0.3$.
To test robustness of the causal effect estimates, we simulated data with two instruments and two exposures (Fig. 1B), with randomly generated non-zero pairwise correlation between the instruments, randomly generated matrix of instrument effect sizes $A$ with determinant greater than $0.05$ , and true causal effects of the exposures on the outcome equal to $c_{X_{1}}=0.2$ and $c_{X_{2}}=0.6$, but estimated causal effects for each exposure separately under the (false) hypothesis that the data was generated from Fig. 1A.
@@
We assumed that a randomly selected subset of at least $K$ eQTLs were causal and all others were associated with the exposures only through LD. For the causal eQTLs, we sampled the vector $\mathbf{a}$ of instrument effect sizes randomly between $0.1$ and $0.3$,
@@
For the simulations on the accuracy of the LD matrix, we generated datasets for the SLC22A3-LPA-PLG locus (centered on chr $6:161089307$) in liver as exposures. Causal effects of the genes on the outcome variable $Y$ were calculated to be $c_{\textit{SLC22A3}}=0.15$, $c_{\textit{LPA}}=-0.05$ and $c_{\textit{PLG}}=-0.27$ when seven instruments were used, and the same values were also used in the other simulations. We compared the situations where $L=3$ (LD threshold of 0.25), $L=4$ (LD threshold of 0.3), $L=5$ (LD threshold of 0.4), $L=6$ (LD threshold of 0.5), and $L=7$ instruments (LD threshold of 0.8) were used. We performed one-sample simulations where the sample size varied between $500-30000$. For the instrument effect sizes, we assumed that three eQTLs were causal for all genes (their effect sizes sampled from a uniform distribution within the range $0.1-0.3$), while additional eQTLs were assumed to be associated with the exposures only through LD.
For the simulations on weak instrument bias, we generated datasets within the same locus and setup as the LD accuracy simulations except here we kept $L=8$ eQTLs (LD threshold of 0.96) and varied the sample size between $500-10000$. For the instrument effect sizes, we again assumed that only three eQTLs were causal for all genes, their effect sizes sampled from a uniform distribution within the range $0.1-0.3$ for strong instruments and within the range $0.001-0.01$ for weak instruments. We confirmed that these instruments were indeed weak/strong by employing the conditional F-statistics computed using the Mendelian Randomization package [30] as well as the MVMR package [31] (S3 Fig).
@@
To simulate the effect of horizontal pleiotropy, we generated datasets for the SLC22A3-LPA-PLG locus with causal effects of the genes on the outcome variable $Y$ set to $c_{\textit{LPA}}=-0.05$ and $c_{\textit{PLG}}=-0.27$ and $c_{\textit{SLC22A3}}\in[0.0,0.4]$. We used eight instruments where only three eQTLs were causal for all genes, their effect sizes sampled from a uniform distribution within the range $0.1-0.3$. We estimated the effects in a misspecified estimation model where we performed MVMR only using LPA-PLG genes, assuming horizontal pleiotropy through one unknown/unmeasured gene SLC22A3 in the same locus.
@@
We compared the least squares (LS) (eq. (7)) and GMM (eq. (10)) estimators to seven other multi-variable MR methods: Transcriptome-Wide MR (TWMR) [19], the mv_multiple estimator from TwoSampleMR package [32], and the mv_mvivw estimator from the Mendelian Randomization package [30], estimators MVMR-Robust, MVMR-Median, and MVMR-Lasso [33] and estimator MVMR-cML [34].
@@
We used eQTL summary statistics from the Stockholm-Tartu Atherosclerosis Reverse Network Engineering Task (STARNET) study, a genetics-of-gene expression study of tissue samples from blood ($n=481$), atherosclerotic-lesion-free internal mammary artery (MAM, $n=552$), atherosclerotic aortic root (AOR, $n=539$), subcutaneous fat (SF, $n=573$), visceral abdominal fat (VAF, $n=531$), skeletal muscle (SKLM, $n=534$), and liver (LIV, $n=576$) obtained during open thorax surgery of 600 coronary artery disease (CAD) patients [26]. eQTL summary statistics from matching tissues in GTEx [12] were used for replication analyses. Significant cis-eQTL associations were defined as in the original STARNET study (FDR $<5\%$ across all tested SNP-gene combinations where the SNP is within 1Mb up or downstream of the center of the gene).
@@
To define locus boundaries and candidate genes of interest for MVMR, we considered genome-wide significant SNPs, all their linked eQTLs within a $0.5Mb$ radius, and all the genes they are associated with in one or more tissues (“eGenes”). Specifically, we first created tissue-specific outcome data with GWAS summary statistics for all SNPs with non-zero (FDR $<5\%$) eQTL effect in the STARNET summary statistics for the tissue of interest. We filtered these eQTLs by their GWAS p-values, retaining only those with p-values below $5\times 10^{-8}$ in the GWAS study. We then iterated through these eQTLs, sorted by GWAS p-value, such that for the first iteration, the first eQTL would be the first lead SNP. We collected all eQTLs within a $0.5Mb$ radius and with non-zero LD with the lead SNP, and then removed the lead SNP as well as its collection of linked eQTLs from the list, repeating the procedure with the next lead SNP. For each lead SNP and its collection of linked eQTLs, we first removed SNPs in perfect LD with the lead SNP and then treated the remaining eQTLs as instrumental variables in one causal diagram containing all genes associated with these eQTLs either in a specific tissue, or across all tissues, depending on the application, to complete the diagram. This procedure ensures that for each locus of interest, we obtain a a closed causal diagram, where “closed” means that no more genes (exposures) can be added to the diagram that have shared cis-eQTL associations with any eQTLs (instruments) either already in the diagram or linked to eQTLs in the diagram.
We computed conditional F-statistics using the Mendelian Randomization [30] and MVMR [31] packages to verify instrument strength.
@@ Results
Fig 2: Causal effect estimation from data simulating instrument correlations, instrument strength, and horizontal pleiotropy. (A) Causal diagram for the simulation of two exposures $X_{1}$ and $X_{2}$ for an outcome $Y$, with two shared instruments $E_{1}$ and $E_{2}$ with pairwise covariance $\alpha$ and matrix of instrument effect sizes $A=[a_{ij}]$. (B) Difference ($\hat{c}-c$) between estimated and true causal effect of $X_{1}$ (true effect size $c=0.2$) on $Y$ across 1,000 independently simulated datasets for a range of sample sizes. (C,D) Distribution of the difference between estimated and true causal effects for $X_{1}$ (left; true effect size $c=0.2$) and $X_{2}$ (right; true effect size $c=0.6$) across 1,000 independently simulated datasets and a range of sample sizes comparing strong ($0.1-0.3$) vs weak ($0.001-0.01$) instruments. (E) False causal diagram used for inference where one causal exposure is missing. (F-G) Difference between estimated and true causal effect of $LPA$ (left, true effect size $c_{\textit{LPA}}=-0.05$) and $PLG$ (right, true effect size $c_{\textit{PLG}}=-0.27$) on $Y$ across 1,000 independently simulated datasets with sample size 2000 across a range of effect sizes of the hidden gene SLC22A3 mediating horizontal pleiotropic effects of the instruments on $Y$, showing both the estimates with the correct model (green) and the estimates with the hidden exposure model (orange). See Methods for simulation details.
@@
Fig 3: Causal effect estimation from data simulating LD matrix sampling bias. The first and second row show a causal diagram for the simulation of overdetermined systems where the number of causal variants was equal to the number of exposures (first row, three exposures; second row, two exposures) and other variants were only correlated with the exposures and outcome by LD. (A, B, C) Comparison between the GMM estimator using an unbiased (B) and biased (C) LD matrix for the causal effect of simulated SLC22A3, for different LD pruning thresholds (corresponding to models with 3, 4, 5, 6 and 7 instruments). (D, E, F) Comparison between the least-squares, GMM, TWMR, MV multiple (TwoSampleMR), and MVIVW (Mendelian Randomization package) estimators using independently sampled datasets of different sizes for exposures and outcomes to emulate a two-sample MVMR setting. (G, H, I) Difference between estimated and true causal effect of $SLC22A3$ (left, true effect size $c_{\textit{SLC22A3}}=0.15$), $LPA$ (center, true effect size $c_{\textit{LPA}}=-0.05$) and $PLG$ (right, true effect size $c_{\textit{PLG}}=-0.27$) on $Y$ showing comparison between the least-squares, GMM, MVMR Robust, MVMR Median, MVMR Lasso and MVMR cML estimators using simulated datasets from causal diagram (A). All distributions plots show the difference between estimated and true causal effects across 1,000 independently sampled datasets. See Methods for simulation details.
@@ Discussion
Extensive simulations confirmed the validity and usefulness of these theoretical results. Most surprisingly perhaps was the finding that the variance in causal effect estimates remained small, even at high correlation values between the instruments. Moreover we saw no benefit in using the computationally more expensive two-stage least squares estimator instead of the simpler least squares estimator – both were unbiased and had identical variance in all our simulations. As expected, horizontal pleiotropy or misspecified LD matrices in the two-sample setting introduce bias in the estimated causal effects, as in all applications of MR, but using conservative simulation parameters, we observed convergence to the correct estimates at sample sizes at the upper end of what is currently available. Of note, in the one-sample setting, that is, when exposures and outcomes are simulated from the same LD structure, misspecified LD matrices (randomly sampled from a distribution centred around the true LD matrix) did not lead to noticeable bias or increased variance.
@@
Throughout this paper, we have defined regulatory pleiotropy as the situation where the same genetic variants are associated to multiple genes in the same locus resulting in multiple paths by which the variants can affect downstream complex traits. Thus, accounting for (measured) regulatory pleiotropy using MVMR is different from methods that account for unmeasured horizontal pleiotropy. However, the fact that regulatory pleiotropy can extend over multiple cell types and tissues and that data from some potentially relevant cell types or tissues will almost always be missing, implies that the two concepts are interrelated: unmeasured regulatory pleiotropy likely is an important source of horizontal pleiotropy.
If data from multiple tissues are available, our results imply that MVMR must be run strictly speaking by including all gene-tissue pairs with cis-eQTL associations in a given GWAS locus as exposures in a single, large causal model in order to limit the risk of unmeasured pleiotropy. However, even though instrument sets can be highly correlated and the STARNET data contains only seven tissues, it was almost never possible to run such large models, as the requirement on the number of causal variants included in or tagged by the instrument set still stands. To overcome this limitation, we see two ways forward.
@@ Acknowledgments
Conditional F-statistic and causal effect estimation on simulated data. The distribution of the conditional F-statistic and causal effect estimates are shown using two effect size ranges for weak and strong instruments: $0.001-0.01$ (weak) and $0.1-0.3$ (strong) in subplots (A, B, C, D), and $0.001-0.03$ (weak) and $0.8-1.5$ (strong) in subplots (E, F, G, H). The results are based on 2,000 simulations of over-determined systems for ADAMTS7 (true effect size $0.15$) where subplots (A, E) show the distribution of the conditional F-statistic using the Mendelian Randomization package and subplots (B, F) use the MVMR package. Subplots (C, G) display the distribution of causal effect estimates for the GMM estimator, while subplots (D, H) show the the distribution of the causal effect estimates for the mvivw estimator (using the Mendelian Randomization package).
@@
S8 Fig
Mean Explained Variance of the First Principal Component as a function of Linkage Disequilibrium (LD) $r$ (Correlation Coefficient). The red dashed line represents the theoretical expectation for the explained variance, calculated as $\frac{1+r}{2}$, which derives from the eigenvalues of the covariance matrix for two standardized variables with correlation $r$. The blue bars show the empirical mean explained variance of PC1 (Principal Component 1) obtained from simulations with a sample size of 2000 and 2000 repetitions.
S9 Fig
Comparison of Type 1 error rate and Power rate. We estimate here the Type 1 error rate for the estimators Least Squares (A) and GMM (C) and Power rate for the estimators Least Squares (B) and GMM (D) from 2,000 independently simulated datasets for a fixed sample size of 2000, using discrete instruments with randomly generated covariances with real LD values from the locus on Chromosome 15:79124475 locus shared by genes ADAMTS7 and CTSH in the MAM tissue.